OpenStudy (help_people):

The function below shows the number of car owners f(t), in thousands, in a city in different years t: f(t) = 0.25t2 − 0.5t + 3.5 The average rate of change of f(t) from t = 2 to t = 6 is ______ thousand owners per year. Answer for Blank 1:

OpenStudy (astrophysics):

\[\text{average rate of change} = \frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (astrophysics):

b = 6, a = 2

OpenStudy (help_people):

f(6)-f(2)/4 ?@Astrophysics

OpenStudy (help_people):

@IrishBoy123

OpenStudy (help_people):

@DanJS

OpenStudy (help_people):

@oldrin.bataku

OpenStudy (danjs):

yes, the average rate of change for f(t) over some interval is the slope of a secant line connecting the two endpoints

OpenStudy (danjs):

The interval is between th epoints (2 , f(2)) and (6 , f(6))

OpenStudy (help_people):

is it 1.5?

OpenStudy (danjs):

not sure

OpenStudy (help_people):

ok well cant you make sure ?

OpenStudy (danjs):

yeah looks right, 3/2 less i messed up in my head

OpenStudy (help_people):

ok! could you make sure this is correct?

OpenStudy (help_people):

one more is

OpenStudy (help_people):

The scatter plot below shows the number of pages scanned (y) in different number of hours (x) by a scanning machine: Plot ordered pairs 0, 500 and 1, 1500 and 2, 2500 and 3, 3500 and 4, 4500 and 5, 5500 Which function best represents the data shown in the scatter plot? y = 100 + 1,000x y = 100 − 1,000x y = 500 − 1,000x y = 500 + 1,000x

OpenStudy (danjs):

you can see each next point has an increase in 1000 every time for the y, and 1 for the x

OpenStudy (danjs):

so it is a line

OpenStudy (help_people):

d? i think the asnwer i s

OpenStudy (danjs):

you can do it a few ways, this is prolly the fastest slope from any 2 points is +1000 pages/hour y-intercept is when time is zero, and is given in the first point, that is the initial number pages scanned

OpenStudy (danjs):

right

OpenStudy (help_people):

YAY im going to trun what i have to do in bye :))))